OpenStudy (karatechopper):

Angel has $10.60 in a jar that contains only quarters and dimes. There are 100 coins in the jar. How many dimes are in Angel's jar?

OpenStudy (karatechopper):

please show your work

OpenStudy (accessdenied):

Let \( q := \text{number of quarters} \) and \( d := \text{number of dimes} \). We can model the number of coins in the jar by the equation: \( q + d = 100 \) The total value of the coins in the jar is given by the equation: \( 0.25q + 0.1d = 10.6 \) Solve for \(q\) and \(d\).

OpenStudy (ash2326):

1 dime= 0.1$ 1 cent= 0.01 $ It's given there are 100 coins in the jar so Let the no. of dimes be x No. of cents be y so \[x+y=100................................(1)\] Also they add up to $10.60 so \[0.1x+0.01y=10.60.................(2)\] Can you solve now @karatechopper

OpenStudy (experimentx):

Oops .. man didn't see that there were 100 coins non negative integral solution of 0.25x+0.1y=10.60 x+y=100

OpenStudy (anonymous):

Try out Wolfram Alpha to solve the 2 equations. http://www.wolframalpha.com/input/?i=10.60+%3D+.25x+%2B+.1y%2C+x+%2B+y+%3D+100

OpenStudy (karatechopper):

ok i did all that but i got way f on my answer..can u show me how to work it?